# Contrastive explainable clustering with differential privacy

### v1

@@ 4.1 NonPrivateApproxFC for $k$-median
The second step. We analyze the problem with modified demands $d^{\prime}$. We will group fractional centers from the solution $(\bar{x},\bar{y})$ to create a new solution $(x^{\prime},y^{\prime})$ with cost at most $2\bar{C}_{LP}$ such that $y^{\prime}_{i}=0$ for each $i\notin N^{\prime}$ and $y^{\prime}_{i}\geq 1/2$ for each $i\in N^{\prime}$. We also ensure that $y^{\prime}_{C}\geq 1/2$ in this step, i.e., $C$ will be a fractional center after this.
The next lemma states that for each $j\in N$, there are at least $1/2$ of the total fractional centers within the distance of $2\bar{C}_{j}$.
Lemma 4.
Proof.
Let $s(j)$ be $j$’s closest neighbor location in $N^{\prime}$, the first term above is independent of $y^{\prime}$ and the minimum value of $y^{\prime}_{j}$ is $1/2$. We now construct a $\{1/2,1\}$-integral solution $(\hat{x},\hat{y})$ with no greater cost. Sort the location $j\in N^{\prime},j\neq C$ in the decreasing order of the weight $d^{\prime}_{j}c_{c_{s}(j)j}$ and put $C$ to the first of the sequence, set $\hat{y}_{j}=1$ for the first $2k-n^{\prime}$ locations and $\hat{y}_{j}=1/2$ for the rest. By doing that, we minimize the cost by assigning heaviest weights $d^{\prime}_{j}c_{c_{s}(j)j}$ to the maximum multiplier (i.e., $1$) while assigning lightest weights $d^{\prime}_{j}c_{c_{s}(j)j}$ to the minimum multiplier (i.e., $1/2$) for each $j\in N^{\prime},j\neq C$. Any feasible $1/2$-restricted solution must have $y^{\prime}_{C}=1$ to satisfy the constraint of $C$ so that the contribution of $\hat{y}_{C}$ is the same as its of $y^{\prime}_{C}$. It follows that the cost of $(\hat{x},\hat{y})$ is no more than the cost of $(x^{\prime},y^{\prime})$. ∎
The third step. This step is similar to the part of Step 3 of Charikar et al. (1999) that converts a $\{1/2,1\}$-integral solution to an integral solution with the cost increases at most by $2$. We note that there are two types of center $\hat{y}_{j}=1/2$ and $\hat{y}=1$, hence there are two different processes. All centers $j$ with $\hat{y}_{j}=1$ are kept while more than half of centers $j$ with $\hat{y}_{j}=1/2$ are removed. Since we show that $\hat{y}_{C}=1$ in the previous step, $C$ is always chosen by this step and hence guarantees the constraint of $C$.
Theorem 3.
For the metric k-median problem, the algorithm above outputs an $8$-approximation solution.
Proof.
It is obvious that the optimal of the LP relaxation is the lower bound of the optimal of the integer program. While constructing an integer solution for the LP relaxation with the modified demands, Lemma 13 incurs a cost of $2\bar{C}_{LP}$ and the third step multiplies the cost by a factor of $2$, making the cost of the solution (to the LP) at most $4\bar{C}_{LP}$. Transforming the integer solution of the modidied demands to a solution of the original input adds an additive cost of $4\bar{C}_{LP}$by Lemma 12 and the Theorem follows. ∎
Corollary 1.
Running peCluster with NonPrivateApproxFC be the above K-median algorithm, with probability at least $1-\beta$, $S^{(i)}_{\epsilon}$ is a $(w,t)$-approximation of $OPT_{i}$–the optimal K-median with a center fixed at position $x_{i}$, in which:
$w$ $=8(1+\alpha)$
4.2 NonPrivateApproxFC for $k$-means
In this section, we describe an algorithm that is a $25$-approximation of the k-means problem with a fixed center. We adapt the work by Kanungo et al. (2002) by adding a fixed center constraint to the single-swap heuristic algorithm. Similar to Kanungo et al. (2002), we need to assume that we are given a discrete set of candidate centers $C$ from which we choose $k$ centers. The optimality is defined in the space of all feasible solutions in $C$, i.e., over all subsets of size $k$ of $C$. We then present how to remove this assumption, with the cost of a small constant additive factor, in the Appendix.
Definition 7.
Let $O=(O_{1},O_{2},\ldots,O_{k})$ be the optimal clustering with $O_{1}$ be the cluster with the fixed center $\sigma$. A set $C\subset\mathbb{R}^{d}$ is a $\gamma$-approximate candidate center set if there exists $\sigma\in\{c_{1},c_{2},\ldots,c_{k}\}\subseteq C$, such that:
$cost(c_{1},c_{2},\ldots,c_{k})$ $\leq(1+\gamma)cost(O),$ (11)
Given $u,v\in\mathbb{R}^{d}$, let $\Delta(u,v)$ denote the squared Euclidean distance between $u,v:\Delta(u,v)=dist^{2}(u,v)$. Let $S\subset\mathbb{R}^{d}$, $\Delta(S,v)=\sum_{u\in S}\Delta(u,v)$. Let $P\subset\mathbb{R}^{d}$, $\Delta_{P}(S)=\sum_{q\in P}\Delta(q,s_{q})$, where $s_{q}$ is the closest point of $S$ to $q$. We drop $P$ when the context is clear.
Let $\sigma$ be the fixed center that must be in the output. Let $C$ be the set of candidate centers, that $\sigma\in C$. We define stability in the context of $k$-means with a fixed center $\sigma$ as follows. We note that it differs from the definition of Kanungo et al. (2002) such that we never swap out the fixed center $\sigma$:
Definition 8.
A set $S$ of k centers that contains the fixed center $\sigma$ is called $1$-stable if: $\Delta\big{(}S\setminus\{s\}\cup\{o\}\big{)}\geq\Delta(S),$ for all $s\in S\setminus\{\sigma\}$, $o\in O\setminus\{\sigma\}$.
Algorithm. We initialize $S^{(0)}$ as a set of $k$ centers form $C$ that $\sigma\in S^{(0)}$. For each set $S^{(i)}$, we perform the swapping iteration:
  • Select one center $s\in S^{(i)}\setminus\sigma$
  • Select one replaced center $s^{\prime}\in C\setminus S^{(i)}$
  • Let $S^{\prime}=S^{(i)}\setminus s\cup s^{\prime}$
Theorem 4.
If $S$ is an $1$-stable $k$-element set of centers, then $\Delta(S)\leq 25\Delta(O)$. Furthermore, if $C$ is a $\frac{\gamma}{25}$-approximate candidate center set, then $S$ is a $(25+\gamma)$-approximate of the actual optimal centroids in the Euclidean space.
Kanungo et al. (2002)’s analysis depends on the fact that all centers in the optimal solution are centroidal, hence they can apply centroidal-Lemma in the Appendix. Our analysis must take into account that $\sigma\in O$ is a fixed center, which is not necessarily centroidal. Therefore, we need to adapt and give special treatment for $\sigma$ as an element of the optimal solution. Also, in the swapping iteration above, we never swap out $\sigma$, since the desired output must always contain $\sigma$, which also alters the swap pairs mapping as in Kanungo et al. (2002).
We adapt the mapping scheme of Kanungo et al. (2002), described in the Appendix with one modification: we set up a fixed capturing pair between $\sigma\in S$ and $\sigma\in O$. It makes sense since $\sigma=o_{\sigma}=s_{\sigma}$, as the closest center in $S$ to $\sigma$ is itself. With it, we always generate a swap partition of $S_{j}$ and $O_{j}$ that both contain $\sigma$. The reason we must have this swap partition (even though we never actually perform this swap on $\sigma$) is that we will sum up the cost occurred by every center in $S$ and $O$ by pair-by-pair from which the total costs $\Delta(S)$ and $\Delta(O)$ are formed (in Lemma 5). For this partition, there are two cases:
  • $|S_{j}|=|O_{j}|=1$, which means the partitions contain only $\sigma$, and hence when they are swapped, nothing changes from $S$ to $S^{\prime}$ and hence do not violate the constraint we set above about not swapping out $\sigma$.
  • $|S_{j}|=|O_{j}|>1$, which means the partitions contain some $o\in O$ that $s_{o}=\sigma$. By the design of the swap pairs mapping, all $s\in S_{j}\setminus\sigma$ will be chosen to be swapped out instead, and $\sigma\in S_{j}$ will never be swapped out.
For each swap pair of $s$ and $o$, we analyze the change in the distortion. For the insertion of $o$, all points in $N_{O}(o)$ are assigned to $o$, leads to the change of distortion to be: $\sum_{q\in N_{O}(o)}\big{(}\Delta(q,o)-\Delta(q,s_{q})\big{)}$.
For all points in $N_{S}(s)$ that lost $s$, they are assigned to new centers, which changes the distortion to be:
$\sum_{q\in N_{S}(s)\setminus N_{O}(o)}$ $\big{(}\Delta(q,s_{o_{q}})-\Delta(q,s)\big{)}\leq\sum_{q\in N_{S}(s)}\big{(}\Delta(q,s_{o_{q}})-\Delta(q,s)\big{)},$
and by the property $3$ of swap pairs, $s_{o_{q}}\neq s$, since $o_{q}\neq o$; the inequality is because for all $q$, $\Delta(q,s_{o_{q}})-\Delta(q,s)\geq 0$.
Lemma 5.
(Full proof in Lemma 15) Let $S$ be $1$-stable set and $O$ be the optimal set of $k$ centers, we have $\Delta(O)-3\Delta(S)+2R\geq 0$, where $R=\sum_{q\in P}\Delta(q,s_{o_{q}})$.
Lemma 6.
(Full proof in Lemma 17) With $R$ and $\alpha$ defined as above: $R\leq 2\Delta(O)+(1+2/\alpha)\Delta(S)$.
Proof of Theorem 4. By the results of Lemma 5 and 6:
00 0 $\leq\Delta(O)-3\Delta(S)+2R$ (12)
$\leq\Delta(O)-3\Delta(S)+4\Delta(O)+2\Delta(S)+\frac{4}{\alpha}\Delta(S)$ (13)
$=5\Delta(O)-(1-4/\alpha)\Delta(S).$ (14)
Using the same argument with Kanungo et al. (2002), we have $5/(1-4/\alpha)\geq\alpha^{2}$ or $(\alpha-5)(\alpha+1)\geq 0$ or $\alpha\leq 5$ which means $\frac{\Delta(S)}{\Delta(O)}=\alpha^{2}\leq 25$, and the Theorem follows.
In Section D.2, we describe how to create a $\gamma$-approximation candidate center set for any constant $\gamma$. Using it to create $C$ as a $\frac{25}{\gamma}$-approximation candidate center set and perform the above algorithm on $C$, the final output $S$ will be a $(25+\gamma)$-approximation.
Corollary 2.
$w$ $=(25+\gamma)(1+\alpha)$
Figure 1: A detailed visualization of our dataset (Albemarle County, Virginia) and analysis includes (a) a Scatter plot of the full dataset with sampled points for contrastive analysis. (b) Comparison of k-median clustering with fixed and non-fixed centroids, both private and non-private. (c) Bar graph showing contrastive explanation differences for differential private and non-private k-median with a fixed centroid. (d) Heatmap indicating k-median error costs based on fixed centroid choices.
5 Experiments
In this part of our study, we focus on understanding the impact of the privacy budget $\epsilon$ on the trade-off between privacy and accuracy. Additionally, we explore how $\epsilon$ influences the quality of differentially private explanations.
To assess the impact of $\epsilon$ on privacy and accuracy, we will employ a set of metrics: 1) PO (Private Optimal) to evaluate the clustering cost using the differentially private algorithm, 2) PC (Private Contrastive) to determine the cost in fixed centroid differentially private clustering, 3) RO (Regular Optimal) to measure clustering cost in regular, non-private data, and 4) RC (Regular Contrastive) to analyze the cost of fixed centroid clustering in non-private data.
Measurements To measure the quality of explanations, we define two key metrics: 1) APE (Average Private Explanation), which calculates the difference between PC and PO to represent the decrease in utility for clustering as an explanatory output for agents, and 2) AE (Average Explanation), which is the difference between RO and RC, serving as a baseline to assess the extent to which differential privacy impacts the explanations. These metrics will enable us to effectively quantify the explanatory power of our approach.
We study the trade-offs between privacy and utility for the above metrics. The $\epsilon$ value balances privacy and utility: lower $\epsilon$ values prioritize privacy but may compromise utility.
Datasets Our study focuses on activity-based population datasets from Charlottesville City and Albemarle County, Virginia, previously used for mobile vaccine clinic deployment Mehrab et al. (2022). The Charlottesville dataset includes 33,000 individuals, and Albemarle County has about 74,000. Due to our algorithm’s intensive computational demands, we sampled 100 points for contrastive explanations.
Computational intensity of our algorithm stems primarily from the k-median algorithm we employ, which formulates the problem using linear programming. When the input size is large, processing times for this specific k-median algorithm can increase significantly. However, the implementation we use is one of the more straightforward approaches to executing k-median with a fixed centroid, and faster algorithms could potentially be applied to alleviate these concerns. For this study, we have opted not to focus on optimizing the computation time but rather to sample the data points to expedite our results. Additionally, we chose to sample 100 points because this number strikes a balance between computational efficiency and the clarity of the results for the reader. Importantly, the number of sampled points does not substantially impact the outcome of our analysis, allowing us to maintain the integrity of our findings while ensuring readability and accessibility for our audience.
Furthermore, our method is data agnostic and does not rely on the s-sparse feature for sample points. It is sufficiently adaptable to handle sample points from any distribution. This flexibility ensures that our approach is robust and versatile, capable of effectively managing diverse datasets regardless of their sparsity or the specific characteristics of their distribution.
Dimensiality It is important to note that our experiments were conducted on a 2D dataset. This choice was driven by the fact that the initial step of our method involves reducing dimensionality, as outlined in DimReduction (Algorithm 1). This process entails scaling down and normalizing the data to produce a lower-dimensional dataset. By implementing this step at the outset, we effectively avoid the challenges associated with managing high-dimensional datasets, thus simplifying the data preparation phase.
Experimental results. In this section, we focus solely on the Albemarle County dataset, with additional analyses and datasets presented in the Appendix. Figure 1(a) displays the dataset, highlighting the original data points with blue dots and the sampled points used for contrastive explanations with orange ’X’ markers. In Figure 1(b), we observe significant trends in the k-median clustering results for both private and non-private settings. Both the Private Optimal (PO) and Private Contrastive (PC) exhibit a downward trend in the private setting, indicating that as the $\epsilon$ budget increases, the error cost decreases, which underscores the trade-off between privacy and accuracy. The trends for non-private settings are depicted as straight lines, reflecting their independence from the privacy budget and are included for comparative purposes.
Our analysis shows a consistent gap between the Private Optimal (PO) and Private Contrastive (PC) across different levels of $\epsilon$, marking a stable contrastive error that does not fluctuate with changes in the privacy budget. This stability is pivotal as it supports our goal of providing dependable explanations to agents, regardless of the $\epsilon$ value, ensuring that our approach adeptly balances privacy considerations with the demand for meaningful explainability.
The distinction between PO and PC arises because PO is derived from implementing the optimal k-median/k-means algorithm, while PC results from enforcing a centroid’s placement at a specific agent’s location, inherently producing a less optimal outcome due to this constraint. Initially, agents are presented with the optimal outcome (PO). When an agent inquires why a centroid is not positioned at their specific location, we offer a contrastive explanation (PC), which is naturally less optimal than the original result as it evaluates the utility change from hypothetically placing a centroid at that specific location. In this scenario, a higher score signifies lower utility, emphasizing the balance between providing personalized explanations and adhering to privacy constraints.
Figure 1(c) displays the average private explanation (APE) and average explanation (AE) with variance across different explanations. The average values of APE across multiple runs are relatively stable, irrespective of the privacy budget, reinforcing the robustness of our methodology against the constraints imposed by differential privacy.
Lastly, the heatmap in Figure 1(d) provides insights into how the relative spatial positioning affects the contrastive explanation. It vividly shows that the greater the distance a fixed centroid is from the distribution center, the larger the contrastive error, further illustrating the intricate dynamics of our clustering approach.
6 Discussion
Our work explores the design of private explanations for clustering, particularly focusing on the k-median and k-means objectives for Euclidean datasets. We formalize this as the PrivEC problem, which gives a contrastive explanation to each agent equal to the increase in utility when a cluster centroid is strategically positioned near them. The accuracy bounds of our algorithm match private clustering methods, even as they offer explanations for every user, within a defined privacy budget.
The related work in this domain has seen the development of algorithms for contrastive explanations, but our contribution stands out by integrating differential privacy guarantees. Our experiments demonstrate the resilience of our approach. Despite the added layer of providing differential private explanations on top of differential private clustering, the quality of our explanations remains uncompromised.
Appendix A Related work: additional details
The analysis presented in Goethals et al. [2022] considers security concerning contrastive explanations. The authors introduced the concept of the ”explanation linkage attack”, a potential vulnerability that arises when employing instance-based strategies to derive contrastive explanations. To address this concern, they put forth the notion of k-anonymous contrastive explanations. Furthermore, the study highlights the intricate balance between transparency, fairness, and privacy when incorporating k-anonymous explanations. As the degree of privacy constraints is heightened, a discernible trade-off comes into play: the quality of explanations and, consequently, transparency are compromised.
There has been a lot of work on private clustering and facility location, starting with  Gupta et al. [2010], which was followed by a lot of work on other clustering problems in different privacy models, e.g., Huang and Liu [2018], Stemmer [2020], Stemmer and Kaplan [2018], Nissim and Stemmer [2018], Feldman et al. [2017]. Gupta et al. [2010] demonstrated that the additive error bound for points in a metric space involves an $O(\Delta k^{2}\log{n}/\epsilon)$ ) term, where $\Delta$ is the space’s diameter. Consequently, all subsequent work, including ours, assumes points are restricted to a unit ball.
Appendix B Additional proofs for PECluster
Theorem 5.
(Full version of Theorem 1) The solution $c$ and $\text{cost}{(S_{\epsilon})}$ released by Algorithm 3 are $\epsilon$-DP. For all clients $i$, the value $\text{cost}(S_{\epsilon})-\text{cost}{S^{(i)}_{\epsilon}}$ released by Algorithm 4 is $\epsilon$-$x_{i}$-exclusive DP.
Proof.
It follows that $\text{cost}(S_{\epsilon})$ is the direct results (using no private information) of $Y$, which is $\epsilon/2$-differentially private coreset. By the post-processing property, $\text{cost}(S_{\epsilon})$ is $\epsilon/2$-DP (which implies $\epsilon$-DP).
$c$ is the output of DimReverse that is $\epsilon/2$-DP w.r.t. the input $(X_{1},X_{2},\ldots,X_{k})$. By composition, $c$ is $\epsilon$-DP since the input $(X_{1},X_{2},\ldots,X_{k})$ is calculated partially from $Y$, which is $\epsilon/2$-DP.
For each explanations $S^{(i)}_{\epsilon}$, let $D,D^{\prime}:D\setminus\{x_{i}\}\sim D^{\prime}\setminus\{x_{i}\}$, let $S^{(i)}_{\epsilon}(Y^{D},x_{i})$ be the value of $S^{(i)}_{\epsilon}$ with input dataset $D$ (and $Y^{D}$ as the private coreset of $D$ respectively) and a fixed center $x_{i}$, let $T=\{Y:S^{(i)}_{\epsilon}(Y,x_{i})\in S\}$, we have $D\sim D^{\prime}$ and therefore:
$\Pr[S^{(i)}_{\epsilon}(D,x_{i})\in S]$ $=\Pr[Y^{D}\in T]$ (15)
$\leq e^{\epsilon/2}\Pr[Y^{D^{\prime}}\in T]$ (16)
$=e^{\epsilon/2}\Pr[S^{(i)}_{\epsilon}(D^{\prime},x_{i})\in T],$ (17)
which implies that $\text{cost}(S^{(i)}_{\epsilon})-\text{cost}(S_{\epsilon})$ is $\epsilon$-$x_{i}$-exclusive DP, since $\text{cost}(S_{\epsilon})$ is $\epsilon/2$-DP (which implies $\epsilon/2$-$x_{i}$-exclusive DP). ∎
Lemma 7.
(Full version of Lemma 1) With probability at least $1-\beta$, $(c_{1},\ldots,c_{k})$ is a $(w,t)$-approximation of $OPT$, where:
$w$ $=w^{\prime\prime}(1+\alpha),$ (18)
$t$ $=\mathcal{O}_{p,\alpha,w^{\prime\prime}}\left(\frac{(k/\beta)^{O_{p,\alpha}(1)}}{\epsilon}.\text{polylog}(n/\beta)\right)$ (19)
Proof.
This result derives from Theorem 48 of Ghazi et al. [2020] (From Eq(20) to Eq(24) of Ghazi et al. [2020]). ∎
Lemma 8.
(Full version of Lemma 2) Fix an $i$. If $OPT_{i}\geq w^{\prime\prime}(1+\alpha)OPT+t^{(i)}$, then with probability at least $1-2\beta$, cost($S_{\epsilon}$) and cost($S^{(i)}_{\epsilon}$) released by Algorithm  3 and  4 satisfies that $cost(S^{(i)}_{\epsilon})>cost(S_{\epsilon}$)
@@ Appendix B Additional proofs for PECluster
(Johnson-Lindenstrauss (JL) Lemma Johnson and Lindenstrauss [1984], Dasgupta and Gupta [2003]) Let $v$ be any $d$-dimensional vector. Let $\mathcal{S}$ denote a random $d^{\prime}$-dimensional subspace of $\mathbb{R}^{d}$ and let $\Pi_{S}$ denote the projection from $\mathbb{R}^{d}$ to $\mathcal{S}$. Then, for any $\zeta\in(0,1)$ we have
$\Pr\left[\|v\|_{2}\approx_{1+\zeta}\sqrt{d/d^{\prime}}\|\Pi_{\mathcal{S}}(v)\|_{2}\right]\geq 1-2\exp\left(-\frac{d^{\prime}\zeta^{2}}{100}\right)$ (24)
@@
Theorem 6.
(Full version of Theorem 2) Fix an $i$. With probability at least $1-\beta$, cost($S^{(i)}_{\epsilon}$) released by Algorithm 4 is a $(w,t)$-approximation of the optimal cost $OPT_{i}$, in which:
$w$ $=w^{\prime}(1+\alpha)$
$t$ $=w^{\prime}\mathcal{O}_{k,p}\left(\frac{2^{O_{p,\alpha}(d)}k^{2}\log^{2}{n}}{\epsilon}.\text{polylog}(\frac{n}{\beta}\right),$
and $OPT_{i}$ is the optimal $(k,p)$-clustering cost with a center is fixed at position $x_{i}$.
@@
Let $\tilde{\mathbf{X}}=(\tilde{x}_{1},\tilde{x}_{2},\ldots,\tilde{x}_{n})$ and $X=(x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{n})$, setting $\alpha^{\prime}=0.1\alpha$, applying Lemma 10, we have:
@@
$\Pr\left[\|x\|>\frac{1}{1+\zeta}\sqrt{d/d^{\prime}}\|x^{\prime}\|\right]$ $\geq 1-2\exp(-d^{\prime}\zeta^{2}/100).$ (27)
@@
$\Pr\left[\|x^{\prime}\|<(1+\zeta)\sqrt{d^{\prime}/d}\right]$ $\geq 1-2\exp(-d^{\prime}\zeta^{2}/100).$ (28)
Setting $\zeta=\sqrt{\frac{log(n/\beta)}{0.01}}-1$, $\Lambda=\frac{1}{1+\zeta}\sqrt{d/d^{\prime}}=\sqrt{\frac{0.01}{\log{(n/\beta)}}.\frac{d}{d^{\prime}}}$, we have:
$\Pr[\|x^{\prime}\|<1/\Lambda]$ $\geq 1-2\exp(-\frac{d^{\prime}\zeta^{2}}{100})$ (29)
@@
By union bound on all $i$, then with probability at least $1-2\beta$, $x^{\prime}_{i}=\Lambda\tilde{x}_{i}$ for all $i$. Since $Y$ is the output of PrivateCoreset with input $X$, then by Theorem 38 of Ghazi et al. [2020], $Y$ is a $(0.1\alpha,t)$-coreset of $X^{\prime}$ (with probability at least $1-\beta$), with $t^{\prime}$ as:
$t^{\prime}=O_{p,\alpha}\left(\frac{2^{O_{p,\alpha}(d^{\prime})}k^{2}\log^{2}{n}}{\epsilon}\log\left(\frac{n}{\beta}\right)+1\right)$ (32)
@@
where in $(a)$ we substitute the value of $t^{\prime}$, and $(b)$ is because $d^{\prime}=O_{\alpha}(p^{4}\log(n/\beta))$, and $(c)$ is because $\beta<1$, and the Lemma follows. ∎

### v2

@@ Appendix B Additional proofs for peCluster
It follows that $\text{cost}(S_{\epsilon})$ is the direct results of $Y$, which is $\epsilon/2$-differentially private coreset. By the post-processing property, $\text{cost}(S_{\epsilon})$ is $\epsilon/2$-DP (which implies $\epsilon$-DP).
The output $c$ of the DimReverse algorithm is $\epsilon/2$-differentially private with respect to the input $(X_{1},X_{2},\ldots,X_{k})$, where $X_{i}$ represents the data points in cluster $i$. The overall process achieves $\epsilon$-differential privacy through composition, as $(X_{1},X_{2},\ldots,X_{k})$ is partially derived from $Y$, which itself is $\epsilon/2$-differentially private.
For each explanations $S^{(i)}_{\epsilon}$, let $X,X^{\prime}:X\setminus\{x_{i}\}\sim X^{\prime}\setminus\{x_{i}\}$, i.e., $X$ and $X^{\prime}$ are any two neighbor datasets that differ at exact one data point that is not agent $i$. Let $S^{(i)}_{\epsilon}(Y^{X})$ be the value of $S^{(i)}_{\epsilon}$ with input dataset $X$ (and $Y^{X}$ as the private coreset of $X$ respectively). Note that we specify the dataset $X$ (and $X^{\prime}$) as a parameter of $S^{(i)}_{\epsilon}$ to highlight the original dataset where the explanation comes from (either $X$ or $X^{\prime}$). Fix any set $S$, let $T=\{Y:S^{(i)}_{\epsilon}(Y)\in S\}$ ,i.e., the set of coresets $Y$ that make $S^{(i)}_{\epsilon}(Y)\in S$. Since $X\sim X^{\prime}$, we have:
$\Pr[S^{(i)}_{\epsilon}(Y^{X})\in S]$ $=\Pr[Y^{X}\in T]$ (7)
$\leq e^{\epsilon/2}\Pr[Y^{X^{\prime}}\in T]$ (8)
$=e^{\epsilon/2}\Pr[S^{(i)}_{\epsilon}(Y^{X^{\prime}})\in S],$ (9)
which implies that $\text{cost}(S^{(i)}_{\epsilon})-\text{cost}(S_{\epsilon})$ is $\epsilon$-$x_{i}$-exclusion DP, since $\text{cost}(S_{\epsilon})$ is $\epsilon/2$-DP (which implies $\epsilon/2$-$x_{i}$-exclusion DP). ∎
See 2
@@
Definition 9.
For $p\geq 1,\alpha>0$, $\lambda_{p,\alpha/2}\overset{def}{=}\frac{1+\alpha/2}{((1+\alpha/2)^{1/p}-1)^{p}}$.
@@
(Johnson-Lindenstrauss (JL) Lemma Johnson and Lindenstrauss [1984], Dasgupta and Gupta [2003]) Let $v$ be any $d$-dimensional vector. Let $\mathcal{S}$ denote a random $d^{\prime}$-dimensional subspace of $\mathbb{R}^{d}$ and let $\Pi_{S}$ denote the projection from $\mathbb{R}^{d}$ to $\mathcal{S}$. Then, for any $\tau\in(0,1)$ we have
$\Pr\left[\|v\|_{2}\approx_{1+\tau}\sqrt{d/d^{\prime}}\|\Pi_{\mathcal{S}}(v)\|_{2}\right]\geq 1-2\exp\left(-\frac{d^{\prime}\tau^{2}}{100}\right)$ (14)
@@
See 3
@@
Let $\tilde{\mathbf{X}}=(\tilde{x}_{1},\tilde{x}_{2},\ldots,\tilde{x}_{n})$, i.e., the projected data after applying the transformation $\Pi_{\mathcal{S}}$. Let $X=(x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{n})$, i.e., $\tilde{\mathbf{X}}$ after being clipped by $1/\Lambda$. By setting $\alpha^{\prime}=0.1\alpha$, and applying Lemma 7, we have:
@@
$\Pr\left[\|x\|>\frac{1}{1+\tau}\sqrt{d/d^{\prime}}\|x^{\prime}\|\right]$ $\geq 1-2\exp(-d^{\prime}\tau^{2}/100).$ (17)
@@
$\Pr\left[\|x^{\prime}\|<(1+\tau)\sqrt{d^{\prime}/d}\right]$ $\geq 1-2\exp(-d^{\prime}\tau^{2}/100).$ (18)
Setting $\tau=\sqrt{\frac{log(n/\beta)}{0.01}}-1$, $\Lambda=\frac{1}{1+\zeta}\sqrt{d/d^{\prime}}=\sqrt{\frac{0.01}{\log{(n/\beta)}}.\frac{d}{d^{\prime}}}$, we have:
$\Pr[\|x^{\prime}\|<1/\Lambda]$ $\geq 1-2\exp(-\frac{d^{\prime}\tau^{2}}{100})$ (19)
@@
By union bound on all $i$, then with probability at least $1-2\beta$, $x^{\prime}_{i}=\Lambda\tilde{x}_{i}$ for all $i$. Since $Y$ is the output of PrivateCoreset with input $X^{\prime}$ and $\zeta$, then by Theorem 38 of Ghazi et al. [2020], $Y$ is a $(0.1\alpha,t)$-coreset of $X^{\prime}$ (with probability at least $1-\beta$), with $\alpha=\frac{(100\zeta)^{p}}{10\lambda_{p,\alpha/2}}$ and $t^{\prime}$ as:
$t^{\prime}=O_{p,\alpha}\left(\frac{2^{O_{p,\alpha}(d^{\prime})}k^{2}\log^{2}{n}}{\epsilon}\log\left(\frac{n}{\beta}\right)+1\right).$ (22)
We note that alternatively, given a target approximation parameter $\alpha$, we can set $\zeta=0.01\left(\frac{\alpha}{10\lambda_{p,\alpha/2}}\right)^{1/p}$.
@@
where in $(a)$ we substitute the value of $t^{\prime}$, and $(b)$ is because $d^{\prime}=O_{\alpha}(p^{4}\log(n/\beta))$, and $(c)$ is because $\beta<1$, and the Lemma follows.